# A note on internality of certain differential systems

### v1

@@ Abstract.
To prove the converse, let $(x,y)$ be a generic solution of the system (C). From the conditions (i) and (ii), we obtain that $C(x)$ contains an element $u$ such that either $u^{\prime}=1$ or $u^{\prime}=cu$ for some nonzero $c\in C$ and that there is an element $v\in C(x,y)\setminus C(x)$ such that $v^{\prime}=dv$ for some $d\in C.$ Now let $(x_{1},y_{1})$ be another generic solution of the system (C). Consider the differential $C$-isomorphism $\phi:C(x,y)\rightarrow C(x_{1},y_{1}),$ where $\phi(x)=x_{1}$ and $\phi(y)=y_{1}.$ Let $c_{2}:=v/\phi(v).$ If $u^{\prime}=1$ then let $c_{1}:=u-\phi(u)$ and if $u^{\prime}=cu$ the let $c_{1}:=u/\phi(u).$ In any event, note that $c_{1},c_{2}$ are constants and that $u,v\in C(x_{1},y_{1},c_{1},c_{2}).$ Since $x,y$ are algebraic over $C(u,v),$ they are also algebraic over $C(x_{1},y_{1},c_{1},c_{2})$. This proves that the system (C) is almost internal to constants. ∎

### v2

@@ Abstract.
To prove the converse, let $(x,y)$ be a generic solution of the system (A). From conditions (i) and (ii), we obtain that $C(x)$ contains an element $s$ such that $s^{\prime}=1$ or $s^{\prime}=cs$ for some nonzero $c\in C$ and that there is an element $t\in C(x,y)\setminus C(x)$ such that $t^{\prime}=bt$ for some $d\in C.$ Note that $s=u$ and $t=y^{m}/v.$ Now let $(x_{1},y_{1})$ be another generic solution of the system (A). Consider the differential $C$-isomorphism $\phi:C(x,y)\rightarrow C(x_{1},y_{1}),$ where $\phi(x)=x_{1}$ and $\phi(y)=y_{1}.$ Let $c_{2}:=t/\phi(t).$ If $s^{\prime}=1$ then let $c_{1}:=s-\phi(s)$ and if $s^{\prime}=cs$ then let $c_{1}:=s/\phi(s).$ In any event, note that $c_{1},c_{2}$ are constants and that $s,t\in C(x_{1},y_{1},c_{1},c_{2}).$ Since $x,y$ are algebraic over $C(s,t),$ they are also algebraic over $C(x_{1},y_{1},c_{1},c_{2})$. This shows that the system (A) is almost internal to the constants. ∎
We conclude this article with the following observations. When $g(x)=x,$ we recover Theorem B of [JM20]. The nonorthogonality of the system (A), under the assumption that $x^{\prime}=f(x)$ is orthogonal to the constants, has been examined in [JJP23, Corollary 5.5]. Interestingly, the condition (ii) of Theorem 5 also appears in their result.
Acknowledgement. We thank the anonymous referee of this article for various helpful suggestions and references.